Minnie is holding a yardstick perpendicular to the ground. The shadow of the yardstick is 27 inches long. Minnie’s shadow is 4 feet 6 inches long. How tall is Minnie?
a.5 feet tall
b.5 feet 6 inches tall
c.40.5 inches tall
d.60 inches tall
e.6 feet tall

Answer by mananth(16946)

the yard stick = 36 inches
shadow = 27 inches
shadow of minnie = 54 inches
let minnie be x inches
..
27/36= 54/x
x=36*54/27
x = 72 inches= 6 feet